OpenStudy (anonymous):

What is the perimeter of a square with an area of 36?

OpenStudy (anonymous):

perimeter of a square is P=4s

OpenStudy (anonymous):

and area is A=s^2

OpenStudy (anonymous):

plug in area 36=x^2 and get x=6

OpenStudy (anonymous):

then plug in to the perimeter formula

OpenStudy (anonymous):

s^2 = 36 s = root of 36 s = 36 area = 4*s = 4*6 = 24

OpenStudy (anonymous):

P=4(6)

OpenStudy (anonymous):

Area of Square = 36 Area = s^2 s=6 (since 6^2=36) perimeter of square = 4s final answer = 24 (since 4*6=24)